OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, +/-6) and asymptotes at y = +/- 3/4 x

OpenStudy (anonymous):

@iambatman @Compassionate @jigglypuff314 anyone????

OpenStudy (anonymous):

um thanks, but I did read my book, I just don't really get it...

jigglypuff314 (jigglypuff314):

I don't know, sorry perhaps @ganeshie8 or @iambatman can help you

OpenStudy (anonymous):

thanks! :)

ganeshie8 (ganeshie8):

see if this helps

OpenStudy (anonymous):

I see the equation I have to use... but how to I get a and b ?

OpenStudy (michele_laino):

@Turner from the data or your texr, I see that we have: \[\frac{ b }{ a }=\frac{ 3 }{ 4 }\] and \[c=6\] Now, in the case of an hyperbola, the relationship between a, b and c, is: \[c ^{2}=a ^{2}+b ^{2}\] From the first equation, we can write: \[b=\frac{ 3 }{ 4 }a\] please insert that relationship for b, into the quadratic equation forc, and you will get a and b. the answer is: \[a=\frac{ 24 }{ 5 }\] and \[b=\frac{ 18 }{ 5 }\] please, check all my statements

OpenStudy (michele_laino):

@Turner the standard form, as you know, is: \[\frac{ x ^{2} }{ a ^{2} }-\frac{ y ^{2} }{ b ^{2} }=1\]

OpenStudy (anonymous):

oh I see! So then I would write it like (x^2 / 4.8) - (y^2 / 3.6) ?

OpenStudy (michele_laino):

better is: \[\frac{ 25x ^{2} }{ 576 }-\frac{ 25y ^{2} }{ 324 }=1\]